Shift report 20010807 16:00-24:00

From: Apache (apache@pii3.brahms.bnl.gov)
Date: Tue Aug 07 2001 - 00:18:53 EDT

  • Next message: Claus O. E. Jorgensen: "Re: New Si and BB dN/dEta at 100 Gev/A !!!"

    ------------------------------------------------------------
    Date: Tue Aug  7 00:18:53 2001, Shift: 16:00-24:00
    Supervisor: RKarabowicz, with: HBoggild
    Log book#7 pages 37 - 40
    
    Shift Summary:
    ------------------------------------------------------------
    1. Ramp for RF Storage Cavity till 2000hrs, at 2045 we were
    granted a two-hours Access to the IR, then still no beam.
    2. No runs taken, except of pedestals. 
    RD and others were in the IR to repair T1 and T2. TPCs seems
    to be all right now, and the voltages on them is not to be
    moved ( more information was sent to nbi by Hans Boggild, 
    and is available in BRAHMS' LogBook#7, pages 39 and 40 ).
    Summary of collisions from run#3928 was made. Since then we 
    gathered about 1.7M collisions ( more information in BRAHMS'
    LogBook#7, page 38 ).
    We had some problems with magnet D2 - there was a stand-by
    fault. Magnet Guys fixed it, and said that sometimes
    it helps for MagnetRemoteControl to push File/Reload.
    3. Besides mentioned above we had no detector problems.
    During Access the magnets were off, afterwards we turned
    them back on to:
    D1 - 676 amps, pol B
    D2 - 555 amps, pol B
    D3 - 588 amps, pol B
    D4 - 534 amps, pol B
    D5 - 350 amps, pol B.
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b30 : Tue Aug 07 2001 - 00:19:31 EDT